You have a group of 40 people. You need to pick 20 in order to form a committee. How many different ways can you do this?
Answer #1

I have a group of 40 people.

From 40 people I can choose 20 in 40C20 ways.

Because, the first choice has 40 possibilities. Our second choice has (40-1)=39 possibilities. Proceeding in this way we gat our 20th choice has (40-19)=21 possibilities.

So total number of ways = 40*39*38*.......*21 = 40P20

Now, here order is not necessary. So total number of ways 40P20 / 20! = 40C20